# Classical scattering and fragmentation of clusters of ions in helical confinement

### v1

@@ II Scattering of ions in helical confinement
$=\sum_{i=1}^{N}\dfrac{m}{2}\dot{\textbf{r}}^{2}(u_{i})+\sum_{\begin{subarray}{c}i,k=1\\ k>i\end{subarray}}^{N}\frac{1}{4\pi\epsilon_{0}}\frac{q^{2}}{|\textbf{r}(u_{i})-\textbf{r}(u_{k})|}$
$=$ $\sum_{i=1}^{N}\xi\dot{u}_{i}^{2}+\sum_{\begin{subarray}{c}i,k=1\\ k>i\end{subarray}}^{N}k\left(2\rho^{2}\left(1-\cos(\Delta_{ik})\right)+\left(\dfrac{h\Delta_{ik}}{2\pi}\right)^{2}\right)^{-\frac{1}{2}},$
@@
where $E_{c}$ is the total kinetic energy at which the clusters collide, and $m$ is the ion mass. The momentum is chosen such, that the total momentum during a collision is conserved, i.e. $N_{A}p_{A}+N_{B}p_{B}=0$. Due to the large number of possible equilibrium configurations, we will - for simplicity - only consider a special subset of these equilibrium configurations where the ions are (insofar as possible) spaced equidistant in the parametric coordinates. Some examples for such configurations are visualized in Fig. 1(c)-(e).

### v2

@@ II Scattering of ions in helical confinement
$=\sum_{i=1}^{N}\dfrac{m}{2}\dot{\textbf{r}}^{2}(u_{i})-\sum_{\begin{subarray}{c}i,k=1\\ k>i\end{subarray}}^{N}\frac{1}{4\pi\epsilon_{0}}\frac{q^{2}}{|\textbf{r}(u_{i})-\textbf{r}(u_{k})|}$
$=$ $\sum_{i=1}^{N}\xi\dot{u}_{i}^{2}-\sum_{\begin{subarray}{c}i,k=1\\ k>i\end{subarray}}^{N}k\left(2\rho^{2}\left(1-\cos(\Delta_{ik})\right)+\left(\dfrac{h\Delta_{ik}}{2\pi}\right)^{2}\right)^{-\frac{1}{2}},$
@@
A final comment regarding the numerical calculations: The ion cluster dynamics can show sensitivity to initial conditions. We verified the convergence of our (finite time) numerical calculations with respect to changes of the time step width. Furthermore, we verified the conservation of the total energy throughout simulations.
@@
where $E_{c}$ is the total kinetic energy at which the clusters collide, and $m$ is the ion mass. The momentum is chosen such that the total momentum during a collision is conserved, i.e. $N_{A}p_{A}+N_{B}p_{B}=0$. Due to the large number of possible equilibrium configurations, we will - for simplicity - only consider a special subset of these equilibrium configurations where the ions are (insofar as possible) spaced equidistant in the parametric coordinates. Some examples for such configurations are visualized in Fig. 1(c)-(e).
@@ V Brief summary and conclusions
Experimental realizations of our setup could be envisioned in the context of cold atoms. Helical trapping potentials have already been realized for neutral atoms [54]. Furthermore, the investigated regime of $<30$ ions is certainly accessible with current trapping techniques for cold ions. For example, linear Paul traps can be used for the controlled capture of small particle numbers, ranging from single ions up to larger ensembles of hundreds or even thousands of ions [55]. Finally, on a more speculative side, the here-observed cluster dynamics are (purely phenomenologically) reminiscent of heavy ion collisions.